# Cooperative Multi-Agent Planning with Adaptive Skill Synthesis

### v1

@@ 4 Methods
COMPASS, illustrated in Figure 1, is a decentralized closed-loop framework for cooperative multi-agent systems that continuously incorporate environmental feedback for strategy refinement. The architecture comprises three core components: 1) a VLM-based closed-loop planner that iteratively perceives, reasons, reflects and acts to adaptively complete tasks (Sec. 4.1); 2) an adaptive skill synthesis mechanism for generating executable codes tailored to proposed sub-tasks (Sec. 4.2); and 3) a structured communication protocol that enables agents to share visible entity information under partial observability (Sec. 4.3).
@@ 4.1 VLM-based Closed-Loop Planner
Inspired by recent advances in cognitive architectures for autonomous systems (Tan et al., 2024), COMPASS implements a sophisticated modular planning framework that emulates key aspects of cognitive decision-making. The planner adopts a modular formulation, utilizing four specialized models: Perception, Task Reasoning, Self-Reflection, and Actor. Each model fulfills a distinct yet interconnected role in the decision-making process. The Perception model processes multi-modal inputs, integrating both visual and textual information to build comprehensive environmental understanding. The Task Reasoning model analyzes the perceived information to decompose complex objectives into manageable sub-tasks, ensuring systematic progress toward the final goal. The Self-Reflection model continuously evaluates task execution and outcome quality, enabling adaptive behavior refinement. The Actor model translates plans into actions by selecting and executing the most appropriate skills from the skill library. We next discuss the various components in detail:
@@
Perception forms the foundation of COMPASS’s decision-making capabilities by enabling robust multi-modal understanding of complex environments. Solving complex real-world tasks often involves data of multiple modalities (Wang et al., 2024a), each contributing unique and complementary information for decision-making. We leverage the VLMs’ ability to fuse and analyze a broader spectrum of data, including text- and image-based environment feedback, to enable agents to sense the surrounding environment. The system’s perception mechanism operates at two levels: direct observation processing and collaborative information synthesis. At the direct level, VLMs process raw inputs to extract meaningful features and relationships from both visual and textual data. At the collaborative level, COMPASS addresses the inherent challenge of partial observability in multi-agent systems through an innovative multi-hop communication protocol (detailed in Sec. 4.3) that enables agents to construct a more holistic understanding of their environment by sharing and aggregating observations. This dual-level perception architecture ensures that each agent maintains both detailed local awareness and broader contextual understanding, essential for effective decision-making in complex cooperative tasks.
Figure 2: Visualization of COMPASS’s dynamic task reasoning process in the StarCraft Multi-Agent Challenge (SMACv2) environment. The figure demonstrates how the VLM-based planner decomposes a complex final goal (”defeat all enemy units”) into a sequence of concrete, executable sub-tasks that adapt to the changing battlefield conditions. This closed-loop task decomposition enables efficient coordination among multiple agents under partial observability, as each sub-task provides clear, actionable objectives that agents can execute while maintaining overall mission alignment.
@@
Task Reasoning enables COMPASS to systematically approach complex cooperative challenges through collective task decomposition. Given a simple general final task in the cooperative multi-agent setting, e.g., ”defeat all enemy units”, in order to complete the task more efficiently, agents are required to decompose it into multiple sub-tasks and figure out the right one to focus on, while considering alignment among others (See Figure. 2). COMPASS harnesses the power of VLMs to analyze high-level task instructions in conjunction with environmental feedback and team member objectives to generate tractable sub-tasks that collectively advance the overall mission. As agents act under stochastic, partially observable environments, the task reasoning model continuously adapts its plans, proposing and refining sub-tasks based on emerging situations and progress assessment. This dynamic approach enables COMPASS to maintain strategic coherence while adjusting tactical decisions in response to changing circumstances.
@@
Actor serves as the critical bridge between high-level reasoning and concrete action execution. Building upon recent advances in code-writing language models for embodied control (Liang et al., 2023; Wang et al., 2024d), the Actor leverages the skill library by first identifying relevant skills for the proposed sub-task, then synthesizes perception and self-reflection inputs to select the optimal skill for execution. This streamlined approach ensures efficient skill selection while maintaining task alignment.
@@
Self-Reflection enables COMPASS to continuously evaluate and refine its decision-making processes through systematic performance analysis. COMPASS instantiates the Self-Reflection model as a VLM which takes a sequence of visual results from the last skill execution with corresponding descriptions as input to assess the quality of the decision produced by the Actor and whether the task was completed. Additionally, we also request the VLM to generate verbal self-reflections to provide valuable feedback on the completion of the task.
@@ 4.2 Adaptive Skill Synthesis
COMPASS employs a dynamic skill library that maintains and evolves a collection of executable behaviors. Each skill is represented as an executable Python function with comprehensive documentation describing its functionality and corresponding embedding that enables semantic retrieval. This skill library undergoes continuous refinement through two complementary mechanisms: incremental synthesis, where new skills are generated and existing ones are refined during task execution, and demonstration-based bootstrapping, which initializes the library with behaviors extracted from expert demonstrations.
@@
Bootstrapping However, developing the skill library from scratch requires extensive interactions with environments, which potentially leads to inefficient learning in the early stages. Inspired by offline MARL approaches (Meng et al., 2023; Zhang et al., 2022; Zhu et al., 2025), which leverage pre-collected datasets to enhance sample efficiency, we leverage MAPPO as the behavior policy to collect experiences, which are recorded as video sequences. The VLMs then analyze these demonstrations through a multi-stage process: first identifying key strategic patterns and behavioral primitives, then translating these patterns into executable Python functions with appropriate documentation. This initialization methodology establishes a foundational set of validated skills, substantially reducing the exploration overhead typically required for discovering effective behaviors. The resulting baseline skill library enables efficient task execution from the onset while maintaining the flexibility to evolve through incremental synthesis.
@@ 4.3 Structured Communication Protocol
To facilitate effective collaboration under partial observability, recent LLM-based multi-agent work (Li et al., 2024a; Zhang et al., 2024) employs conversational framework with unconstrained communication protocol. However, while natural language offers flexibility, unrestricted communication can lead to potential hallucinations caused by ambiguous or irrelevant messages between agents. Drawing from advances in structured communication frameworks (Hong et al., 2024) and entity-based MARL (Iqbal et al., 2021; Ding et al., 2023), COMPASS implements a hierarchical communication protocol that focuses on efficient entity-based information sharing and multi-hop propagation. Each agent maintains an observation buffer containing information about entities in its field of view. At each timestep, agents share their local observations, which are then aggregated into a global entity memory accessible to all. COMPASS employs a multi-hop communication mechanism to propagate information about distant entities, enabling agents to build a more holistic observation of the environment by leveraging the collective knowledge of the team.
@@ 5 Experiments
We conducted a comprehensive experimental evaluation of COMPASS to assess its performance and capabilities in complex multi-agent scenarios. Our evaluation focused on the improved StarCraft Multi-Agent Challenge (SMACv2) (Ellis et al., 2024), which provides an ideal testbed for examining cooperative behavior under partial observability and stochasticity. Through systematic experimentation, we investigated two fundamental questions: (1) How does COMPASS perform compared to state-of-the-art MARL methods? (2) What are the individual contributions of the adaptive skill synthesis mechanism? Experiments utilize both open-source (Qwen2-VL-72B) and closed-source VLMs (GPT-4o-mini, Claude-3-Haiku), with Jina AI embeddings for skill retrieval. All results are averaged over 5 seeds to account for environmental stochasticity.
@@ 5.1 Experimental Setup
Scenarios Our evaluation scenarios span three distinct race matchups (Protoss, Terran, and Zerg) and two categories (symmetric and asymmetric), as detailed in Table 1. The symmetric scenarios (5v5) test coordination in balanced engagements, while asymmetric scenarios (5v6) evaluate adaptation to numerical disadvantages. Each race combination presents unique tactical challenges due to different unit abilities and constraints.
Table 1: SMACv2 task scenarios
Task Scenarios Categories
Protoss protoss 5 vs 5 symmetric
protoss 5 vs 6 asymmetric
Terran terran 5 vs 5 symmetric
terran 5 vs 6 asymmetric
Zerg zerg 5 vs 5 symmetric
zerg 5 vs 6 asymmetric
@@
  • Value-Based Methods: QMIX (Rashid et al., 2020) uses a mixing network architecture to decompose joint action-values while maintaining monotonicity constraints.
  • Policy Gradient Methods: MAPPO (Yu et al., 2022) extends PPO to multi-agent settings with the CTDE paradigm. HAPPO (Kuba et al., 2022) performs sequential policy updates by utilizing other agents’ newest policy under the CTDE framework and provably obtains the monotonic policy improvement guarantee. HASAC (Liu et al., 2024) combines the maximum entropy framework with trust region optimization to enhance exploration and coordination.
@@ 5.2 Main Results
Table 2: Comparative performance of COMPASS (with three VLM variants) and state-of-the-art MARL baselines on SMACv2. Median win rates (%) and standard deviations (subscripts) are reported across Protoss, Terran, and Zerg scenarios in symmetric (5v5) and asymmetric (5v6) categories. Results are averaged over 5 seeds. Bold values denote the best performance in each scenario.

### v2

@@ 4 Methods
COMPASS, illustrated in Figure 1, is a decentralized closed-loop framework for cooperative multi-agent systems that continuously incorporate environmental feedback for strategy refinement. The architecture comprises three core components: 1) a VLM-based closed-loop planner that iteratively perceives, reasons, reflects and acts to adaptively complete tasks (Sec. 4.1); 2) an adaptive skill synthesis mechanism for generating executable codes tailored to proposed sub-tasks (Sec. 4.2); and 3) a structured communication protocol that enables agents to share visible entity information under partial observability (Sec. 4.3). The pseudo-code of COMPASS is shown in Appendix.
Figure 2: Visualization of COMPASS’s dynamic task reasoning process in the StarCraft Multi-Agent Challenge (SMACv2) environment. The figure demonstrates how the VLM-based planner decomposes a complex final goal ("defeat all enemy units") into a sequence of concrete, executable sub-tasks that adapt to the changing battlefield conditions. This closed-loop task decomposition enables efficient coordination among multiple agents under partial observability, as each sub-task provides clear, actionable objectives that agents can execute while maintaining overall mission alignment.
@@ 4.1 VLM-based Closed-Loop Planner
Inspired by recent advances in cognitive architectures for autonomous systems Tan et al. (2024), COMPASS implements a sophisticated modular planning framework that emulates key aspects of cognitive decision-making. The planner adopts a modular formulation, utilizing four specialized models: Perception, Task Reasoning, Self-Reflection, and Actor. Each model fulfills a distinct yet interconnected role in the decision-making process. The Perception model processes multi-modal inputs, integrating both visual and textual information to build comprehensive environmental understanding. The Task Reasoning model analyzes the perceived information to decompose complex objectives into manageable sub-tasks, ensuring systematic progress toward the final goal. The Self-Reflection model continuously evaluates task execution and outcome quality, enabling adaptive behavior refinement. The Actor model translates plans into actions by selecting and executing the most appropriate skills from the skill library. We next discuss the various components in detail:
@@
Perception forms the foundation of COMPASS’s decision-making capabilities by enabling robust multi-modal understanding of complex environments. Solving complex real-world tasks often involves data of multiple modalities Wang et al. (2024a), each contributing unique and complementary information for decision-making. We leverage the VLMs’ ability to fuse and analyze a broader spectrum of data, including text- and image-based environment feedback, to enable agents to sense the surrounding environment. The system’s perception mechanism operates at two levels: direct observation processing and collaborative information synthesis. At the direct level, VLMs process raw inputs to extract meaningful features and relationships from both visual and textual data. At the collaborative level, COMPASS addresses the inherent challenge of partial observability in multi-agent systems through an innovative multi-hop communication protocol (detailed in Sec. 4.3) that enables agents to construct a more holistic understanding of their environment by sharing and aggregating observations. This dual-level perception architecture ensures that each agent maintains both detailed local awareness and broader contextual understanding, essential for effective decision-making in complex cooperative tasks.
Figure 3: Illustration of self-reflection. Following skill execution (e.g., ’stalker_cover’) and feedback, COMPASS assesses performance. This analysis guides further skill generation to refine tactics (e.g., registering an ’Enhanced Stalker Script Type Aggressive’).
@@
Task Reasoning enables COMPASS to systematically approach complex cooperative challenges through collective task decomposition. Given a simple general final task in the cooperative multi-agent setting, e.g., "defeat all enemy units", in order to complete the task more efficiently, agents are required to decompose it into multiple sub-tasks and figure out the right one to focus on, while considering alignment among others (See Figure 2). COMPASS harnesses the power of VLMs to analyze high-level task instructions in conjunction with environmental feedback and team member objectives to generate tractable sub-tasks that collectively advance the overall mission. As agents act under stochastic, partially observable environments, the task reasoning model continuously adapts its plans, proposing and refining sub-tasks based on emerging situations and progress assessment. This dynamic approach enables COMPASS to maintain strategic coherence while adjusting tactical decisions in response to changing circumstances.
@@
Actor serves as the critical bridge between high-level reasoning and concrete action execution. Building upon recent advances in code-writing language models for embodied control Liang et al. (2023); Wang et al. (2024d), the Actor leverages the skill library by first identifying relevant skills for the proposed sub-task, then synthesizes perception and self-reflection inputs to select the optimal skill for execution. This streamlined approach ensures efficient skill selection while maintaining task alignment.
@@
Self-Reflection enables COMPASS to continuously evaluate and refine its decision-making processes through systematic performance analysis (See Figure 3). COMPASS instantiates the Self-Reflection model as a VLM which takes a sequence of visual results from the last skill execution with corresponding descriptions as input to assess the quality of the decision produced by the Actor and whether the task was completed. Additionally, we also request the VLM to generate verbal self-reflections to provide valuable feedback on the completion of the task.
Figure 4: Overview of Adaptive Skill Synthesis. VLMs perform (Top) Bootstrapping by analyzing offline data for initial Tactic Analysis and Skill Generation into a Skill Library. (Bottom) Incremental Synthesis uses Task Reasoning to dynamically generate or enhance code-based skills, evolving the library for new tasks. The skills follow a structured decision-making pipeline with two core components: score_target(unit) for dynamic target prioritization and control_logic() for coordinating behavior. Textual observations are parsed into structured data (obs_data), mapping raw text to attributes, e.g., "Can move North: yes" -> can_move=’north’: True.
@@ 4.2 Adaptive Skill Synthesis
COMPASS employs a dynamic skill library that maintains and evolves a collection of executable behaviors. Each skill is represented as an executable Python function with comprehensive documentation describing its functionality and corresponding embedding that enables semantic retrieval. This skill library undergoes continuous refinement through two complementary mechanisms (Figure 4): incremental synthesis, where new skills are generated and existing ones are refined during task execution, and demonstration-based bootstrapping, which initializes the library with behaviors extracted from expert demonstrations. For further details and visualizations of example synthesized tactical behaviors, please refer to the Skill Analysis section in Appendix.
@@
Bootstrapping However, developing the skill library from scratch requires extensive interactions with environments, which potentially leads to inefficient learning in the early stages. Inspired by offline MARL approaches Meng et al. (2023); Zhang et al. (2022); Zhu et al. (2025), which leverage pre-collected datasets to enhance sample efficiency, we leverage MAPPO as the behavior policy to collect experiences, which are recorded as video sequences. The VLMs then analyze these demonstrations through a multi-stage process: first identifying key strategic patterns and behavioral primitives, then translating these patterns into executable Python functions with appropriate documentation. This initialization methodology establishes a foundational set of validated skills, substantially reducing the exploration overhead typically required for discovering effective behaviors. The resulting baseline skill library enables efficient task execution from the onset while maintaining the flexibility to evolve through incremental synthesis.
@@ 4.3 Structured Communication Protocol
To facilitate effective collaboration under partial observability, recent LLM-based multi-agent work Li et al. (2024a); Zhang et al. (2024) employs conversational framework with unconstrained communication protocol. However, while natural language offers flexibility, unrestricted communication can lead to potential hallucinations caused by ambiguous or irrelevant messages between agents. Drawing from advances in structured communication frameworks Hong et al. (2024) and entity-based MARL Iqbal et al. (2021); Ding et al. (2023), COMPASS implements a hierarchical communication protocol that focuses on efficient entity-based information sharing and multi-hop propagation (Figure 5). Each agent maintains an observation buffer containing information about entities in its field of view. At each timestep, agents share their local observations, which are then aggregated into a global entity memory accessible to all. COMPASS employs a multi-hop communication mechanism to propagate information about distant entities, enabling agents to build a more holistic observation of the environment by leveraging the collective knowledge of the team.
@@ 5 Experiments
We conducted a comprehensive experimental evaluation of COMPASS to assess its performance and capabilities in complex multi-agent scenarios. Our evaluation focused on the improved StarCraft Multi-Agent Challenge (SMACv2) Ellis et al. (2024), which provides an ideal testbed for examining cooperative behavior under partial observability and stochasticity. Through systematic experimentation, we investigated two fundamental questions: (1) How does COMPASS perform compared to state-of-the-art MARL methods? (2) What are the individual contributions of each component in COMPASS? Experiments utilize both open-source (Qwen2-VL-72B) and closed-source VLMs (GPT-4o-mini, Claude-3-Haiku), with Jina AI embeddings for skill retrieval. All results are averaged over 5 seeds to account for environmental stochasticity. Token usage is approximately 0.4 million per episode.
@@ 5.1 Experimental Setup
Scenarios Our evaluation scenarios span three distinct race matchups (Protoss, Terran, and Zerg) and two categories (symmetric and asymmetric), as detailed in Appendix. The symmetric scenarios (5v5) test coordination in balanced engagements, while asymmetric scenarios (5v6) evaluate adaptation to numerical disadvantages. Each race combination presents unique tactical challenges due to different unit abilities and constraints. We followed the setting p=0 in the SMACv2 original paper (i.e., prob_obs_enemy: 0.0 in the .yaml file), meaning that only the first agent to initially spot a specific enemy unit can continue observing it, introducing the Extended Partial Observability Challenge, which baselines struggled with.
@@
  • Value-Based Methods: QMIX Rashid et al. (2020) uses a mixing network architecture to decompose joint action-values while maintaining monotonicity constraints.
  • Policy Gradient Methods: MAPPO Yu et al. (2022) extends PPO to multi-agent settings with the CTDE paradigm. HAPPO Kuba et al. (2022) performs sequential policy updates by utilizing other agents’ newest policy under the CTDE framework and provably obtains the monotonic policy improvement guarantee. HASAC Liu et al. (2024) combines the maximum entropy framework with trust region optimization to enhance exploration and coordination.
@@ 5.2 Main Results
Table 1: Comparative performance of COMPASS (with three VLM variants: G-4o=GPT-4o-mini, C-Hk=Claude-3-Haiku, Q2-VL=Qwen2-VL-72B) and state-of-the-art MARL baselines on SMACv2. Median win rates (%) and standard deviations (subscripts) are reported across Protoss, Terran, and Zerg scenarios in symmetric (5v5) and asymmetric (5v6) categories. Results are averaged over 5 seeds. Bold values denote the best performance in each scenario.
@@
Moreover, COMPASS demonstrates particular advantages in scenarios with sparse reward, where traditional MARL approaches significantly underperform (near-zero win rates).
5.3 Ablation Studies
PROTOSS TERRAN ZERG
5V5 $0.35_{0.06}$ $0.24_{0.04}$ $0.06_{0.01}$
5V6 $0.04_{0.05}$ $0.06_{0.02}$ $0.02_{0.03}$
Skill Initialization To evaluate the impact of our skill initialization, we analyze the performance of COMPASS using only the initialized skill library derived from expert demonstrations. The results in Table 2 demonstrate that skill initialization alone achieves non-trivial performance across different scenarios, particularly in symmetric matchups. Moreover, the gap between initialized skills and COMPASS underscores the necessity of incremental skill synthesis. A script example for skill initialization is in Appendix.
Communication To demonstrate the critical role of communication, we evaluated COMPASS on Protoss 5v5 under the Extended Partial Observability setting, using only local information without multi-hop propagation. The resulting win rate with GPT-4o-mini decreased to $0.06_{0.04}$, a significant drop from $0.57$ with full communication. This degradation occurs because the Extended Partial Observability setting restricts direct enemy visibility to the first agent that initially spots it. The VLMs generated control logic heavily relies on the presence of enemies in the local observation to determine engagement and targeting. Without communication relaying enemy positions, agents other than the discoverer cannot ’see’ enemies known to teammates, even if within attack range. Consequently, their control logic frequently defaults to ’no enemy’ behaviors, such as moving towards allies or executing random default actions, preventing effective target engagement and coordinated attacks, thus drastically reducing combat effectiveness and the overall win rate.
Self Reflection In order to show the effectiveness of self-reflection, we evaluate the performance of COMPASS w/o self-reflection on protoss 5 vs 5. Removing the module leads to a drop -10% ($0.47_{0.04}$).
Visual information We tested visual information contribution by omitting the image inputs. Intuitively, this challenges the agents to rely solely on uni-modal textual information, potentially leading to a loss of spatial understanding. Empirical results show a 10% drop in performance when visuals are omitted. Comparing the VLM outputs shows that without visual information, understanding map boundaries relies on interpreting textual cues (e.g., ’West (unavailable movement)’), which might be derived indirectly from action availability rather than direct observation. With visual input, the VLM can directly perceive these crucial spatial details like map boundaries from the image itself. This lack of spatial awareness can result in less informed tactical decisions regarding movement and positioning.
6 Conclusion
Appendix A Pseudocode
The pseudo-code of the COMPASS algorithm is shown in Pseudocode 1.
Initialize:
skill_manager.bootstrap(demonstration_data)
agent_state $\leftarrow$ environment.reset(agent_id)
local_memory $\leftarrow$ initialize_local_memory()
global_memory $\leftarrow$ initialize_global_memory()
previous_action_result $\leftarrow$ None
while True do
       // 1. Communication Phase
      local_observations $\leftarrow$ agent_state.get_observations()
       communication_protocol.share_local_observations(agent_id, local_observations, global_memory)
       global_entity_info $\leftarrow$ communication_protocol.get_global_memory.update(global_memory)
      // 2. Perception Phase
      processed_state $\leftarrow$ vlm_perception.process( raw_observation=agent_state, communication_data=global_entity_info, local_memory=local_memory )
       local_memory.update(processed_state)
      // 3. Self-Reflection Phase
      if previous_action_result $\neq$ None then
             reflection_feedback $\leftarrow$ vlm_self_reflection.reflect( previous_action_result )
      // 4. Task Reasoning Phase
      sub_task $\leftarrow$ vlm_task_reasoning.props_subtask( processed_state, overall_goal, reflection_feedback )
      // 5. Skill Generation Phase
      new_skill_code $\leftarrow$ vlm_skill_generator.generate_skill(sub_task, processed_state)
       skill_manager.add_skill(new_skill_code, sub_task)
      // 6. Actor Phase
      relevant_skills $\leftarrow$ skill_manager.retrieve_skills(sub_task)
       chosen_skill_code $\leftarrow$ vlm_actor.select_skill( sub_task, relevant_skills, processed_state )
      // 7. Execution Phase
      (next_agent_state, reward, done, info) $\leftarrow$ environment.step( agent_id, chosen_skill_code )
      agent_state $\leftarrow$ next_agent_state
       previous_action_result $\leftarrow$ (chosen_skill_code, info, reward, done)
       local_memory.add_action(chosen_skill_code)
      if done then
             break
Algorithm 1 COMPASS Agent Decision-Making Loop
Appendix B Implementation Details
COMPASS integrates VLMs to process multi-modal inputs and generate executable skills in two stages. Each skill follows a standardized interface:
1def skill_template(obs: str):
2 obs_data = parse_obs(obs)
3 def score_target(unit):
4 ...
5 return score
6 def control_logic():
7 ...
8 return atomic_action
Listing 1: Interface of Generated skills.
The skills follow a structured decision-making pipeline with two core components: score_target(unit) and control_logic():
  • score_target(unit): Dynamically calculates a threat/priority score by evaluating unit type, health, distance, formations, and matchups to guide optimal attack/heal targeting decisions.
  • control_logic(): Dynamically coordinates unit behavior by integrating observations, target priorities, and pathfinding to execute role-optimized strategies (e.g., stalkers attack colossus at max range while moving away from zealots).
COMPASS evolves skills through iterative refinement and task-guided synthesis:
  • iterative refinement: When errors occur during skill execution, VLMs analyze the error messages and attempt to fix the bugs.
  • task-guided synthesis: When a new task is proposed, VLMs first determine whether a new skill needs to be generated to align with the task. If necessary, VLMs generate new score_target or control_logic components to fulfill the task requirements and integrate them with the existing code to construct a new skill.
For example, if the task is: "Implement an aggressive advance movement pattern for the colossus unit when enemy stalkers are within sight range and allies are positioned to provide covering fire," and the current skills in the library are not aggressive enough, the VLMs will refine the control_logic to implement a more aggressive behavior pattern.
COMPASS maintains a global entity memory, where agents act as nodes connected if within sight range. Each node stores information about visible allies and enemies. When Agent A observes entities, it propagates updates through adjacent nodes recursively, up to max_hops (default = 3). This allows agents to infer off-screen threats via ally intermediaries. For implementation details, see ./common/memory/global_memory.py.
Appendix C Skill Analysis
We now analyze COMPASS’s capability to synthesize and execute diverse tactical behaviors. COMPASS develops four key tactical patterns: (1) An exponentially-scaled focus fire implementation that coordinates multiple units’ target selection based on allied attacker density (Figure 6), (2) A position-aware kiting mechanism that maintains optimal engagement ranges while managing unit positioning relative to threats (Figure 7), (3) A formation-based isolation tactic that enables systematic target elimination through coordinated unit movements (Figure 8), and (4) An area-of-effect (AOE) tactic that maximizes splash damage through cluster density calculation (Figure 9). These synthesized skills exhibit clear strategic intent while maintaining interpretability.
Appendix D Baseline training results
For QMIX, we utilized Epymarl [^3] https://github.com/uoe-agents/epymarl, and for others, we used HARL [^4] https://github.com/PKU-MARL/HARL. Task difficulty was set to 5v5 for SYMMETRIC tasks and 5v6 for ASYMMETRIC tasks; the 5v6 scenario introduces a significant force disadvantage (20% outnumbered vs 10% in 10v11).
Figure 10: Baseline training results on SMACv2.
Appendix E Environment Settings and More Results
Table 3: SMACv2 task scenarios
Task Scenarios Categories
Protoss protoss 5 vs 5 symmetric
protoss 5 vs 6 asymmetric
Terran terran 5 vs 5 symmetric
terran 5 vs 6 asymmetric
Zerg zerg 5 vs 5 symmetric
zerg 5 vs 6 asymmetric